# Asymptotic evaluation of the information processing capacity in reservoir computing

### v1

@@ Abstract
The squared error normalized by the target output is known as the information processing capacity (IPC) and is used to evaluate the performance of reservoir computing (RC). Since RC aims to learn the relationship between input and output time series, we should evaluate the IPC for infinitely long data rather than the IPC for finite-length data. To evaluate the IPC for infinitely long data using the IPC for finite-length data, we use an asymptotic expansion of the IPC and the least-squares method. Then, we show the validity of our method by numerical simulations.
@@ 1 Introduction
Since many kinds of data, e.g. video, audio, and stock prices, have time correlation, machine learning of time-series data is an important issue. Recurrent neural networks (RNNs) can store past input by recursively connecting hidden nodes [1] and can approximate the relationship between input and output time series with arbitrary accuracy [2]. Backpropagation through time (BPTT) is mainly used to train RNNs, but it is difficult to optimize network parameters due to the gradient vanishing or the gradient explosion [3]. Many variants of RNNs, such as LSTM [4] and GRU [5], have been proposed to solve the difficulty of training and have been very successful. However, BPTT calculations become slower for longer training data.
@@
An echo state network (ESN) [6] is a kind of RNNs, which can finish training quickly by fixing the recurrent connections at the initial value and optimizing only the linear transformation of the readout layer. Not limited to neural networks, a linear combination of nonlinear dynamical systems can be used to approximate the relationship between input and output time series and is called a reservoir computing (RC) system [7]. RC systems can also approximate the relationship between input and output time series with arbitrary accuracy [8, 9]. Since no optimization is performed other than the linear transformation, an RC system is often inferior in performance to LSTM and other methods. However, it has the advantage that training finishes quickly by only calculating the pseudoinverse matrix. Therefore, it is more convenient than LSTM and other methods in situations where the target to be optimized changes frequently.
@@
The performance of an RC system is evaluated by the mean squared error or the squared error normalized by the target output, and the latter is called the information processing capacity (IPC) [10, 11]. The IPC ranges between 00 and $1$, representing the approximation accuracy. Since the goal of RC is to learn the relationship between input and output time series, it is necessary to evaluate the IPC for infinitely long data, not the IPC for finite-length data. The simplest method of estimating the IPC for infinitely long data is to calculate the IPC for sufficiently long data. However, their difference remains and should be removed as much as possible. In this paper, we estimate the IPC for infinitely long data using the asymptotic expansion of the IPC and the least-squares method.
@@
This paper is organized as follows. Section 2 reviews RC and its performance index, the IPC. In Section 3, we derive the asymptotic form of the IPC, and in Section 4, we show the validity of our method by numerical simulations. Section 5 summarizes the paper.
@@ 2 Review of RC and the IPC
Here, $u_{t}\in\mathbb{R}^{d_{0}},x_{t}\in\mathbb{R}^{d_{1}},y_{t}\in\mathbb{R}^{d_{2}}$ are the values of the input, the hidden nodes, and the reservoir output at time $t$, respectively. To simplify the notation, we rewrite $(x_{t}^{\top},1)^{\top}$ as $x_{t}$, and $(w^{\top},b)^{\top}$ as $w$, and Eq. (2) is expressed by
@@
To obtain $x_{t}$ from Eq. (1), the values of the hidden nodes at a certain time $-\tau\,(<t)$ are required as the initial value. The initial value is not optimized in RC. Instead, we employ $f$, which reduces the dependence on the initial value through the recursive equation Eq. (1), and take a sufficiently large $\tau$. Furthermore, in most cases, $f$ reduces the dependence on the past input. For example, the hidden node values of the ESN, a typical example of RC systems, are given by [12]
@@
The cost function for an RC system is usually given by the mean squared error between the reservoir output sequence $(y_{1},\ldots,y_{T})$ and the target output sequence $(\hat{y}_{1},\ldots,\hat{y}_{T})$ for the input sequence $(\ldots,u_{1},\ldots,u_{T})$,  [^1] Although not used in this paper, a regularization term for $w$ may be added to the cost function.
@@
In addition to the mean squared error, a quantity called the IPC, defined below,
$1-\frac{\min_{w}\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}-w^{\top}x_{t}\|^{2}}{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}\|^{2}},$ (7)
@@
is also used as a performance index for RC. Due to the normalization by the target output, the IPC ranges from 00 to $1$, representing the accuracy of the approximation. Since the purpose of RC is to learn the relationship between input and output time series from a finite-length input/output set (training data), the actual performance is given by Eq. (7) with $T\to\infty$. Therefore, we should estimate the IPC for infinitely long training data from the IPC for finite-length training data. The simplest estimation method is to use sufficiently long training data. However, this method has the problem that we cannot evaluate the deviation from the limit value.
3 Asymptotic expansion of the IPC
@@
In this section, we estimate the IPC for infinitely long data using the asymptotic expansion. First, we summarize the RC dynamics by referring to Ref. [10, 8, 9]. From Eq. (1), the values of the hidden nodes of the RC system at time $t$ are determined by the initial value $x_{-\tau}$ and the input sequence $(u_{-\tau},\ldots,u_{t})$. As we have explained, the RC system is usually designed to reduce dependence on the initial value and the past input. Hence, taking $\tau\to\infty$, we can consider that $x_{t}$ is determined only by the input sequence $(u_{t-s})_{s=0}^{\infty}$. Due to time-independence of Eq. (1), if we give the same input sequence at two different times, $(u_{t-s})_{s=0}^{\infty}=(v_{t^{\prime}-s})_{s=0}^{\infty}$, the hidden node values at $t$ with the input sequence $(u_{t-s})_{s=0}^{\infty}$ and those at $t^{\prime}$ with $(v_{t^{\prime}-s})_{s=0}^{\infty}$ are the same. This means that the values of the hidden nodes are determined by a sequence of real numbers, and the hidden nodes of the RC system can be considered as the mapping from a real number sequence (input sequence) to a real vector (hidden node variables). To simplify the notation, in the following, the input sequence $(u_{t-s})_{s=0}^{\infty}$ is represented as $u^{\prime}_{t}$, and ${\cal U}\subseteq(\mathbb{R}^{d_{0}})^{\mathbb{N}}$ denotes the set of input sequences. We express the target output function as $\hat{y}\colon{\cal U}\to\mathbb{R}^{d_{2}}$ and the reservoir function as $x\colon{\cal U}\to\mathbb{R}^{d_{1}+1}$. We assume the components of $x$, namely, $x_{1},\ldots,x_{d},x_{d+1}$, are linearly independent.
@@
Next, we consider the stochasticity of the input sequence. Following Ref. [10, 9], we assume that $U=(U_{t})_{t\in\mathbb{Z}}$ is a stationary ergodic process. We use the notation $U^{\prime}_{t}=(U_{s})_{s\leq t}$, and in particular $U^{\prime}=(U_{t})_{t\leq 0}$. The target output sequence is given by $\{\hat{y}(U^{\prime}_{1}),...,\hat{y}(U^{\prime}_{T}),...,\hat{y}(U^{\prime}_{T+T^{\prime}})\}$, and we divide it into training data $\{(U^{\prime}_{t},\hat{y}(U^{\prime}_{t}))\}_{t=1}^{T}$ and test data $\{(U^{\prime}_{t},\hat{y}(U^{\prime}_{t}))\}_{t=T+1}^{T+T^{\prime}}$. We assume $T^{\prime}=O(T)$ in this paper. The readout layer is optimized by the training data,
$w_{T}=$ $\min_{w}\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (8)
@@
where we use the notation $\|x(U^{\prime}_{t})\|^{2}=x(U^{\prime}_{t})^{\top}x(U^{\prime}_{t})$. The IPCs for the training and the test data at $w_{T}$ is called the training IPC and the test IPC, respectively, which are given by
$C_{T}=$ $1-\frac{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})-w_{T}^{\top}x(U^{\prime}_{t})\|^{2}}{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})\|^{2}}=1-\frac{l_{T}(w_{T})}{\mu_{T}},$ (9)
$C^{\prime}_{T,T^{\prime}}=$ $1-\frac{\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\|\hat{y}(U^{\prime}_{t})-w_{T}^{\top}x(U^{\prime}_{t})\|^{2}}{\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\|\hat{y}(U^{\prime}_{t})\|^{2}}=1-\frac{l^{\prime}_{T^{\prime}}(w_{T})}{\mu_{T^{\prime}}}.$ (10)
@@
$l_{T}(w)=$ $\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (11)
$l^{\prime}_{T^{\prime}}(w)=$ $\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (12)
$\mu_{T}=$ $\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})\|^{2}.$ (13)
Now, we derive the asymptotic forms of the training and the test IPC. In this paper, the IPC for infinitely long training data is called the true IPC $C_{0}$, and $w$ at that time is called the true parameter $w_{0}$. Since RC and the optimized value $w_{T}$ can be considered a linear regression model and its optimal parameter, the asymptotic theory of linear regression models in Ref. [13] can be applied. First, we introduce the notation,
$l(w,U^{\prime}_{t})=$ $\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (14)
$\mu(U^{\prime}_{t})=$ $\|\hat{y}(U^{\prime}_{t})\|^{2},$ (15)
and assume
$\mathbb{E}[l(w,U^{\prime})]<\infty,\ \ \mathbb{E}[\mu(U^{\prime})]<\infty.$ (16)
Then, from the law of large numbers, $l_{T}(w)$ converges in probability to $l(w)=\mathbb{E}[l(w,U^{\prime})]$ for $T\to\infty$, and $\mu_{T}$ to $\mu_{0}=\mathbb{E}[\mu(U^{\prime})]$. We find that $C_{0}$ and $w_{0}$ can be expressed by
@@
Next, we introduce $\zeta(U^{\prime}_{t})=\mu(U^{\prime}_{t})-\mu_{0}$ and define the following quantities,
$I_{\infty}(w)=$ $\frac{1}{4}\sum_{t=0}^{\infty}{\rm Cov}(\nabla_{w}l(w,U^{\prime}_{0}),\nabla_{w}l(w,U^{\prime}_{t})^{\top})+\frac{1}{4}\sum_{t=1}^{\infty}{\rm Cov}(\nabla_{w}l(w,U^{\prime}_{0}),\nabla_{w}l(w,U^{\prime}_{-t})^{\top}),$ (19)
$J=$ $\frac{1}{2}\frac{\partial^{2}l(w)}{\partial w\partial w}\biggr{|}_{w=w_{0}},\ \ J_{ij,kl}=\mathbb{E}[x_{i}(U^{\prime})\,x_{k}(U^{\prime})\,\delta_{jl}],$ (20)
$V_{\zeta,\infty}=$ $\sum_{t=0}^{\infty}{\rm Cov}(\zeta(U^{\prime}_{0}),\zeta(U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(\zeta(U^{\prime}_{0}),\zeta(U^{\prime}_{-t})),$ (21)
$C_{l,\zeta}(w)=$ $\sum_{t=0}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0})\,,\zeta(U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0})\,,\zeta(U^{\prime}_{-t})),$ (22)
We further assume that $I_{\infty}(w),J,V_{\zeta,\infty},C_{l,\zeta}$(w), and $V_{l,\infty}(w)$ are finite matrices. Below, we express $I_{\infty}(w_{0}),C_{l,\zeta}(w_{0})$, and $V_{l,\infty}(w_{0})$ as $I_{\infty},C_{l,\zeta}$, and $V_{l,\infty}$, respectively. Then, the following quantities,
$\xi_{T}=$ $\frac{1}{2}J^{-1/2}\nabla_{w}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}(l(w,U^{\prime}_{t})-l(w))\Bigr{|}_{w=w_{0}}=\frac{1}{2\sqrt{T}}J^{-1/2}\sum_{t=1}^{T}\nabla_{w}l(w_{0},U^{\prime}_{t})\,,$ (24)
$\xi^{\prime}_{T^{\prime}}=$ $\frac{1}{2}J^{-1/2}\nabla_{w}\frac{1}{\sqrt{T^{\prime}}}\sum_{t=T+1}^{T+T^{\prime}}(l(w,U^{\prime}_{t})-l(w))\Bigr{|}_{w=w_{0}}=\frac{1}{2\sqrt{T^{\prime}}}J^{-1/2}\sum_{t=T+1}^{T+T^{\prime}}\nabla_{w}l(w_{0},U^{\prime}_{t})\,,$ (25)

### v2

@@ Abstract
Reservoir computing (RC) is becoming increasingly important because of its short training time. The squared error normalized by the target output is called the information processing capacity (IPC) and is used to evaluate the performance of an RC system. Since RC aims to learn the relationship between input and output time series, we should evaluate the IPC for infinitely long data rather than the IPC for finite-length data. However, a method for estimating it has not been established. We evaluated the IPC for infinitely long data using the asymptotic expansion of the IPC and weighted least-squares fitting. Then, we showed the validity of our method by numerical simulations. This work makes the performance evaluation of RC more evident.
@@ 1 Introduction
Since many kinds of data, e.g. video, audio, and stock prices, have time correlation, machine learning of time-series data is an important issue. Recurrent neural networks (RNNs) can store past input by recursively connecting hidden nodes [rumelhart1986learning] and can approximate the relationship between input and output time series with arbitrary accuracy [schafer2006recurrent]. Backpropagation through time (BPTT) is mainly used to train RNNs, but it is difficult to optimize network parameters due to the gradient vanishing or the gradient explosion [bengio1994learning]. Many variants of RNNs, such as LSTM [hochreiter1997long] and GRU [cho2014learning], have been proposed to solve the difficulty of training and have been very successful. However, training with BPTT takes a long time.
@@
An echo state network (ESN) [jaeger2004harnessing] is a kind of RNNs, which can finish training quickly by fixing the recurrent connections at the initial value and optimizing only the linear transformation of the readout layer. Not limited to neural networks, a linear combination of nonlinear dynamical systems can be used to approximate the relationship between input and output time series and is called a reservoir computing (RC) system [tanaka2019recent]. RC systems can also approximate the relationship between input and output time series with arbitrary accuracy [grigoryeva2018echo, gonon2019reservoir]. Since no optimization is performed other than the linear transformation, an RC system is often inferior in performance to LSTM and other methods. However, RC has the advantage that training is finished quickly by simply calculating the pseudoinverse matrix. It has different demands than large-scale neural networks, which require large amounts of data and high computational costs. Recently, RC has been the subject of active research [yan2024emerging], including improvements in reservoir dynamics [palacios2024role, peng2024reservoir], physical implementation [gao2024toward, lee2024task], and theoretical analysis [ballarin2025memory].
@@
The performance of an RC system is evaluated by the mean squared error (MSE) or the squared error normalized by the target output, and the latter is called the information processing capacity (IPC) [dambre2012information, kubota2021unifying]. Through normalization, the IPC takes a value between 0 and $1$, making it an intuitive index of approximation accuracy. Since the goal of RC is to learn the relationship between input and output time series, it is necessary to evaluate the IPC for infinitely long data, not the IPC for finite-length data. Previous studies have been careful not to overestimate the IPCs that vanish in the infinite time limit, but those that do not vanish in the infinite time limit are substituted with the IPCs for sufficiently long data [dambre2012information, kubota2021unifying, vettelschoss2021information, schulte2023refined, takasu2025neuronal]. However, the difference between the IPC for finite and infinite data lengths still exists and should be removed as much as possible. In this paper, we estimate the IPC for infinitely long data using the asymptotic expansion of the IPC and the weighted least squares method. We also propose a robust method to determine whether the IPC becomes zero in the infinite time limit.
@@
This paper is organized as follows. Section 2 reviews RC and its performance index, the IPC. In Section 3, we derive the asymptotic form of the IPC and explain the method for estimating the IPC. In Section 4, we show the validity of our method by numerical simulations. Section 5 summarizes the paper.
@@ 2 Review of RC and the IPC
Here, $u_{t}\in\mathbb{R}^{d_{0}},x_{t}\in\mathbb{R}^{d_{1}},y_{t}\in\mathbb{R}^{d_{2}}$ are the values of the input, the hidden nodes, and the reservoir output at time $t$, respectively. To simplify the notation, we rewrite $(x_{t}^{\top},1)^{\top}$ as $x_{t}$, $(w^{\top},b)^{\top}$ as $w$, and Eq. (2) can be rewritten as
@@
To obtain $x_{t}$ from Eq. (1), the values of the hidden nodes at a certain time $\tau\,(<t)$ are required as the initial value. The initial value is not optimized in RC. Instead, we employ $f$, which reduces the dependence on the past input through the recursive equation Eq. (1), and take a sufficiently large $\tau$ to decrease the initial value dependence. For example, the hidden node values of the ESN, a typical example of RC systems, are given by [jaeger2002tutorial]
@@
The cost function for an RC system is usually given by the MSE between the reservoir output sequence $(y_{1},\ldots,y_{T})$ and the target output sequence $(\hat{y}_{1},\ldots,\hat{y}_{T})$ for the input sequence $(\ldots,u_{1},\ldots,u_{T})$,  [^1] Although not used in this paper, a regularization term for $w$ may be added to the cost function.
@@
In addition to the MSE, a quantity called the IPC, defined below,
$C_{T}=1-\frac{\min_{w}\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}-w^{\top}x_{t}\|^{2}}{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}\|^{2}},$ (7)
@@
is also used as a performance index for RC. Due to the normalization by the target output, the IPC ranges from 0 to $1$, representing the accuracy of the approximation. Since the purpose of RC is to learn the relationship between input and output time series from a finite-length input/output set (training data), the actual performance is given by $\lim_{T\to\infty}C_{T}$. Therefore, we should estimate the IPC for infinitely long training data from the IPC for finite-length training data. We call $C_{0}\equiv\lim_{T\to\infty}C_{T}$ the true IPC to distinguish it from the training IPC for finite-length data. The simplest estimation method is to use sufficiently long training data. Furthermore, previous studies have introduced a threshold to Eq. (7) to avoid overestimating the IPC, which takes 0 in infinite time limit [dambre2012information, kubota2021unifying, vettelschoss2021information, schulte2023refined, takasu2025neuronal]. In this paper, we call their IPC the empirical IPC defined by
$C_{T}^{\rm emp}=\theta(C_{T}-C_{T,{\rm th}})\,C_{T},$ (8)
where $\theta$ is the Heaviside step function and $C_{T,{\rm th}}$ is the threshold determined by $\chi^{2}$ distribution [dambre2012information] or the random shuffle surrogate [kubota2021unifying]. This method has the problem that we cannot evaluate the deviation from the limit value when the IPC for infinite data length is nonzero.
3 Evaluation of the IPC by the asymptotic expansion
3.1 Asymptotic expansion of the IPC
@@
In this section, we estimate the IPC for infinitely long data using the asymptotic expansion. First, we summarize the RC dynamics by referring to Refs. [dambre2012information, grigoryeva2018echo, gonon2019reservoir]. From Eq. (1), the values of the hidden nodes of the RC system at time $t$ are determined by the initial value $x_{-\tau}$ and the input sequence $(u_{-\tau},\ldots,u_{t})$. As we have explained, the RC system is usually designed to reduce dependence on the initial value and the past input. Hence, taking $\tau\to\infty$, we can consider that $x_{t}$ is determined only by the input sequence $(u_{t-s})_{s=0}^{\infty}$. Due to time-independence of Eq. (1), if we give the same input sequence at two different times, $(u_{t-s})_{s=0}^{\infty}=(v_{t^{\prime}-s})_{s=0}^{\infty}$, the hidden node values at $t$ with the input sequence $(u_{t-s})_{s=0}^{\infty}$ and those at $t^{\prime}$ with $(v_{t^{\prime}-s})_{s=0}^{\infty}$ are the same. This means that the values of the hidden nodes are determined by a sequence of real numbers, and the hidden nodes of the RC system can be considered as the mapping from a real number sequence (input sequence) to a real vector (hidden node variables). To simplify the notation, in the following, the input sequence $(u_{t-s})_{s=0}^{\infty}$ is represented as $u^{\prime}_{t}$, and ${\cal U}\subseteq(\mathbb{R}^{d_{0}})^{\mathbb{N}}$ denotes the set of input sequences. We express the target output function as $\hat{y}\colon{\cal U}\to\mathbb{R}^{d_{2}}$ and the reservoir function as $x\colon{\cal U}\to\mathbb{R}^{d_{1}+1}$. We assume that the components of $x$, that is, $x_{1},\ldots,x_{d_{1}},x_{d_{1}+1}$, are linearly independent.
@@
Next, we consider the stochasticity of the input sequence following Ref. [gonon2019reservoir]. Let $U=(U_{t})_{t\in\mathbb{Z}}$ be a stochastic process. We use the notation $U^{\prime}_{t}=(U_{s})_{s\leq t}$, and in particular $U^{\prime}=(U_{t})_{t\leq 0}$. Since $\hat{y}$ and $x$ are measurable functions in general, the composite functions, $\hat{y}\circ U^{\prime}_{t}$ and $x\circ U^{\prime}_{t}$, are random variables. We divide the target output sequence $\{\hat{y}\circ U^{\prime}_{1},...,\hat{y}\circ U^{\prime}_{T},...,\hat{y}\circ U^{\prime}_{T+T^{\prime}}\}$ into training data $\{(U^{\prime}_{t},\hat{y}\circ U^{\prime}_{t})\}_{t=1}^{T}$ and test data $\{(U^{\prime}_{t},\hat{y}\circ U^{\prime}_{t})\}_{t=T+1}^{T+T^{\prime}}$. We assume $T^{\prime}=O(T)$ in this paper. The readout layer is optimized by the training data,
$w_{T}(u^{\prime}_{T})=$ $\min_{w}\frac{1}{T}\sum_{t=1}^{T}(\hat{y}(u^{\prime}_{t})-w^{\top}x(u^{\prime}_{t}))^{2},$ (9)
@@
where we use the notation $(\hat{y}(u^{\prime}_{t}))^{2}=\hat{y}(u^{\prime}_{t})^{\top}\hat{y}(u^{\prime}_{t})$. The IPCs for the training and test data at $w_{T}$ are called the training IPC and the test IPC, respectively, which are given by
$C_{T}(u^{\prime}_{T})=$ $1-\frac{\frac{1}{T}\sum_{t=1}^{T}(\hat{y}(u^{\prime}_{t})-w_{T}(u^{\prime}_{T})^{\top}x(u^{\prime}_{t}))^{2}}{\frac{1}{T}\sum_{t=1}^{T}(\hat{y}(u^{\prime}_{t}))^{2}}$
$=$ $1-\frac{\bar{l}_{T}(w_{T}(u^{\prime}_{T}),u^{\prime}_{T})}{\bar{\mu}_{T}(u^{\prime}_{T})},$ (10)
$C^{\prime}_{T,T^{\prime}}(u^{\prime}_{T+T^{\prime}})=$ $1-\frac{\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}(\hat{y}(u^{\prime}_{t})-w_{T}(u^{\prime}_{T})^{\top}x(u^{\prime}_{t}))^{2}}{\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}(\hat{y}(u^{\prime}_{t}))^{2}}$
$=$ $1-\frac{\bar{l}^{\prime}_{T,T^{\prime}}(w_{T}(u^{\prime}_{T}),u^{\prime}_{T+T^{\prime}})}{\bar{\mu}_{T^{\prime}}(u^{\prime}_{T+T^{\prime}})}.$ (11)
@@
$\bar{l}_{T}(w,u^{\prime}_{T})=$ $\frac{1}{T}\sum_{t=1}^{T}(\hat{y}(u^{\prime}_{t})-w^{\top}x(u^{\prime}_{t}))^{2},$ (12)
$\bar{l}^{\prime}_{T,T^{\prime}}(w,u^{\prime}_{T+T^{\prime}})=$ $\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}(\hat{y}(u^{\prime}_{t})-w^{\top}x(u^{\prime}_{t}))^{2},$ (13)
$\bar{\mu}_{T}(u^{\prime}_{T})=$ $\frac{1}{T}\sum_{t=1}^{T}(\hat{y}(u^{\prime}_{t}))^{2}.$ (14)
Now, we derive the asymptotic forms of the training and the test IPC. As defined in Section 2, the IPC for infinitely long training data is called the true IPC $C_{0}$, and $w$ at that time is called the true parameter $w_{0}$. Since RC and the optimized value $w_{T}$ can be considered a linear regression model and its optimal parameter, the asymptotic theory of linear regression models in Ref. [white2014asymptotic] can be applied. First, we introduce the notation,
$l(w,u^{\prime}_{t})=$ $(\hat{y}(u^{\prime}_{t})-w^{\top}x(u^{\prime}_{t}))^{2},$ (15)
$\mu(u^{\prime}_{t})=$ $\hat{y}(u^{\prime}_{t})^{2},$ (16)
and define the following quantities;
$I_{\infty}(w)=$ $\frac{1}{4}\sum_{t=0}^{\infty}{\rm Cov}(\nabla_{w}l(w,U^{\prime}_{0}),\nabla_{w}l(w,U^{\prime}_{t})^{\top})$
$+\frac{1}{4}\sum_{t=1}^{\infty}{\rm Cov}(\nabla_{w}l(w,U^{\prime}_{0}),\nabla_{w}l(w,U^{\prime}_{-t})^{\top}),$ (17)
$J=$ $\frac{1}{2}\frac{\partial^{2}\mathbb{E}[l(w,U^{\prime})]}{\partial w\partial w}\biggr|_{w=w_{0}},\ \ J_{ij,kl}=\mathbb{E}[x_{i}(U^{\prime})\,x_{k}(U^{\prime})\,\delta_{jl}],$ (18)
$V_{\mu,\infty}=$ $\sum_{t=0}^{\infty}{\rm Cov}(\mu(U^{\prime}_{0}),\mu(U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(\mu(U^{\prime}_{0}),\mu(U^{\prime}_{-t})),$ (19)
$C_{l,\mu}(w)=$ $\sum_{t=0}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0}),\mu(U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0}),\mu(U^{\prime}_{-t})),$ (20)
We assume that the matrices above are finite. We also assume that the central limit theorem (CLT) holds for $l(w,U^{\prime}_{t})$, $\nabla l(w,U^{\prime}_{t})$, and $\mu(U^{\prime}_{t})$,
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}(l(w,U^{\prime}_{t})-\mathbb{E}[l(w,U^{\prime})])\xrightarrow{d}$ ${\cal N}(0,V_{l,\infty}(w)),$ (22)
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}(\nabla_{w}l(w,U^{\prime}_{t})-\nabla_{w}\mathbb{E}[l(w,U^{\prime})])\xrightarrow{d}$ ${\cal N}(0,I_{\infty}(w)),$ (23)
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}(\mu(U^{\prime}_{t})-\mathbb{E}[\mu(U^{\prime})])\xrightarrow{d}$ ${\cal N}(0,V_{\mu,\infty}),$ (24)
and the law of large numbers (LLN) holds for $x(U^{\prime}_{t})x(U^{\prime}_{t})^{\top}$,
$\frac{1}{T}\sum_{t=1}^{T}x(U^{\prime}_{t})x(U^{\prime}_{t})^{\top}\xrightarrow{p}$ $J.$ (25)
Using the notation $l(w)\equiv\mathbb{E}[l(w,U^{\prime})]$ and $\mu_{0}\equiv\mathbb{E}[\mu(U^{\prime})]$, we find that $C_{0}$ and $w_{0}$ can be expressed by
@@
In the following, we express $I_{\infty}(w_{0}),C_{l,\mu}(w_{0})$, and $V_{l,\infty}(w_{0})$ as $I_{\infty},C_{l,\mu}$, and $V_{l,\infty}$, respectively. From Eq. (23), the following quantities,
$\xi_{T}=$ $\frac{1}{2}J^{-1/2}\nabla_{w}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}(l(w,U^{\prime}_{t})-l(w))\Bigr|_{w=w_{0}}$
$=$ $\frac{1}{2\sqrt{T}}J^{-1/2}\sum_{t=1}^{T}\nabla_{w}l(w_{0},U^{\prime}_{t})\,$, (28)
$\xi^{\prime}_{T^{\prime}}=$ $\frac{1}{2}J^{-1/2}\nabla_{w}\frac{1}{\sqrt{T^{\prime}}}\sum_{t=T+1}^{T+T^{\prime}}(l(w,U^{\prime}_{t})-l(w))\Bigr|_{w=w_{0}}$
$=$ $\frac{1}{2\sqrt{T^{\prime}}}J^{-1/2}\sum_{t=T+1}^{T+T^{\prime}}\nabla_{w}l(w_{0},U^{\prime}_{t})\,$, (29)